It takes you half an hour to travel 26 miles to work. What is you average speed? What formula did you use?
Answer by mananth(16949)

speed = distance / time
distance = 26 miles
time = 1/2 hour
speed = 26 / (1/20
speed = 26*2
speed = 52 mph
...
m.ananth@hotmail.ca